ARDUINO Hi I'm trying to write a code for a personal project, but I haven't used...

ARDUINO

Hi I'm trying to write a code for a personal project, but I haven't used much arduino before. It's supposed to be a function where it takes in motor speed as an int and motor direction as a bool. The PWM is 8 bits. So it's supposed to be an asynchronous function where the blue is always at 50%, then it gets more green as it has a higher speed forward and more red as it has a higher speed backwards. It's supposed to be async, so running constantly without a delay. There should be if and for loops to find values. Thank you!
Answer #1

n= no. of bits

f= frequency

s=slip

p=no. of pole

f1=n*f

ns=120*f1/p

nr1=ns*(1-s)

nr2=ns*(2+s)

if s>.5

print('blue')

print('nr1')

if s<.5

print('green')

print('nr2')

end

end